A boy is inside a bus moving at 5.0m/s toward the east. He is walking at 1.0 m/s toward the front. What is the velocity of the boy as observed by (a) a passenger sitting inside the bus and (b) another passenger waiting for the bus at a nearby bus stop?
1
Expert's answer
2019-08-20T10:10:43-0400

For relative velocities it is necessary to establish:


"V_{P\/A}=V_{P\/B}+V_{B\/A}"



Where:

  • Velocity ​​of the particle (child) with respect to A (Observer passenger)

"V_{P\/A}"

  • Bus velocity (reference B) with respect to A (observer passenger)

"V_{B\/A}"

  • Particle velocity P (child) with respect to B (buss)

"V_{p\/B}"



Now.

Part A

Velocity of the child with respect to the bus:

"V_{P/B}=1.0\frac{m}{s}"


Velocity of the bus with respect to the passenger. (moves at the same velocity as the bus)

"V_{B/A}=0\frac{m}{s}"


Now, evaluating numerically

"V_{P/A}=1.0\frac{m}{s}+0\frac{m}{s}"


The velocity relative to the child on the bus is:

"V_{p/A}=1.0\frac{m}{s}"


Part B

Velocity of the child with respect to the bus.

"V_{P/B}=1.0\frac{m}{s}"


velocity of the bus with respect to the observer.

"V_{B/A}=5.0\frac{m}{s}"


Numerically evaluating


"V_{P/A}=5.0\frac{m}{s}+1.0\frac{m}{s}"


The relative speed of the child with respect to waiting outside the bus is:


"V_{P/A}=6.0\frac{m}{s}"